OpenStudy (anonymous):

1 divided by the cubed root of X to the negative 6 power.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

1/x^(-6/3)=1/x^(-2) that is equal to x^2

OpenStudy (anonymous):

laws used \[\sqrt[n]{x}^{a}=x^{a/n}\] and \[x^{-a}=1/x^{a}\]

OpenStudy (anonymous):

\[1\div (X ^{3})^{-6} \] Does your problem look like this?

OpenStudy (anonymous):

1 divided by the 3rd root of x . X is under the radical and has a negative 6 exponent.

OpenStudy (anonymous):